生命テクノサイエンス演習 2 単位 (必修)

Seminar in Biological Technology  生命テクノサイエンスコース教員

【授業目的】修士論文のテーマに関する演習を行い研究能力を高める． 【授業概要】所属する研究室において，修士論文のテーマに関する演習を行い， 発表，討論することを通じて，データ処理能力，研究能力を高めるとともに， 産業や医療の現場において必要な知識とスキルを習得する．本科目は，工業 に関する科目である． 【授業形式】演習 【キーワード】研究に関する討論，口頭発表 【先行科目】 [先行科目] 【関連科目】『生命テクノサイエンス研究論文』 (0.5)，『生命テクノサイエンス 特別実験』 (0.5)，『生命テクノサイエンス論文輪講』 (0.5) 【履修要件】なし 【履修上の注意】なし 【到達目標】研究方法の習得 【授業計画】 1. 研究発表 1 と議論 2. 研究発表 2 と議論 3. 研究発表 3 と議論 4. 研究発表 4 と議論 5. 研究発表 5 と議論 6. 研究発表 6 と議論 7. 研究発表 7 と議論 8. 研究発表 8 と議論 9. 研究発表 9 と議論 10. 研究発表 10 と議論 11. 研究発表 11 と議論 12. 研究発表 12 と議論 13. 研究発表 13 と議論 14. 研究発表 14 と議論 15. 研究発表 15 と議論 16. 研究発表 16 と議論 【成績評価基準】修士論文を評価 【教科書】なし【参考書】研究論文【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216741 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 生物事務室(M703) 【備考】各々 1 時間の予習と復習が，各 2 時間の授業の理解と単位取得のために

必要である.

Seminar in Biological Technology 2 units (compulsory)

Teacher of course

Target⟩ The goal of this lecture is to aquire the knowledge necessary for research

related to the Master’s thesis. Outline⟩ Student elevates the reaseach ability and skill in the individual laboratory

through presentation and active discussion as to the subject of study. In addition, the research skills in industrial and medical application reseaches should be acquired.

Style⟩ Excercise �������⟩ discussion for research work, oral presentation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “生命テクノサイエンス研究論文”(0.5), “Biological science

and technology laboratory”(0.5), “Practice for undrstanding scientific papers in bioogical technology”(0.5)

Requirement⟩ N/A Notice⟩ N/A Goal⟩ To learn how to perform research

Schedule⟩

1. Research presentation 1 and discussion

2. Research presentation 2 and discussion

3. Research presentation 3 and discussion

4. Research presentation 4 and discussion

5. Research presentation 5 and discussion

6. Research presentation 6 and discussion

7. Research presentation 7 and discussion

8. Research presentation 8 and discussion

9. Research presentation 9 and discussion

10. Research presentation 10 and discussion

11. Research presentation 11 and discussion

12. Research presentation 12 and discussion

13. Research presentation 13 and discussion

14. Research presentation 14 and discussion

15. Research presentation 15 and discussion

16. Research presentation 16 and discussion Evaluation Criteria⟩ Evaluate master thesis Textbook⟩ N/A Reference⟩ Research papers

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216741

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Administration Office(M703)

Note⟩ Preparation and review of 1 hour each are required for understanding of every class with two hours and taking the credits.